Circular No. 7373

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 2000M IN NGC 6389 A. Dimai, Cortina d’Ampezzo, Italy, reports the discovery of an ap-

parent supernova (mag 16.5–17.0) by Marco Migliardi from unfiltered CCD images taken on Feb. 27.17 and 27.20 UT during the Col Druscie Remote Observatory Supernova Search with the 0.5-m reflector. A third image showing the new object was obtained by Migliardi on Feb. 28.186. H. Yamaoka, Kyushu University, reports his measurement of the position of SN 2000M from Migliardi’s discovery image: α = 17^{h}32^{m}39^{s}.61, δ = +16^{o}23′54′′.9 (equinox 2000.0); Migliardi reports that the new object is 3′′ west and 13′′ south of the nucleus of NGC 6389. An unfiltered image taken on 1999 July 24 shows nothing at the position of SN 2000M (limiting mag ∼ 19.0), and it does not appear on the Palomar Sky Survey. W. Li, University of California at Berkeley, writes that he obtained a confirmation image of SN 2000M at mag ∼ 16.8 (unfiltered) with the Katzman Auotmatic Imaging Telescope (KAIT) on Mar. 1.5 UT. The position end figures of the new object are 39^{s}.67, 54′′.1, being offset 1′′.6 west and 12′′.6 south from the nucleus of NGC 6389. A KAIT image of the field taken on 1999 Sept. 11 showed nothing at the position of SN 2000M (limiting mag ∼ 19.5).

SUPERNOVA 2000L IN UGC 5520 S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian Center for

Astrophysics, report that a spectrum of SN 2000L, obtained by M. Calkins on Mar. 1.3 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), reveals it to be a type-II supernova. The spectrum exhibits a blue continuum with P-Cyg Balmer lines. Narrow superimposed Hα emission yields a recession velocity of 3370 km/s for the host galaxy. The supernova expansion velocity, measured to the trough of the Hβ line profile, is 5300 km/s.

COMET C/1999 S4 (LINEAR) Visual m_{1} estimates: 1999 Dec. 30.34 UT, 14.3 (N. Biver, Oahu, HI,

0.26-m reflector); 2000 Jan. 8.75, 14.6 (W. Hasubick, Buchloe, Germany, 0.44-m reflector); Feb. 2.80, 13.5 (R. J. Bouma, Groningen, The Netherlands, 0.25-m reflector); 29.24, 13.4 (M. Linnolt, Haiku, HI, 0.20-m refl.).

2000 March 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green